OpenStudy (anonymous):

1. Simplify. -4^2-(-3)^2 2. Simplify. 7-3(5-9)^2 divided by 4 multiplied by (-2) 3. 2/3 - 1/3 (1/2-5/6)^2 divided by 8/9 4. Use the distributive property to complete the statement. 3(2x-7y)=6x

OpenStudy (anonymous):

@MARC_ Show work

Nnesha (nnesha):

(-3)^2 is same as -3 times -3 which is equal to = ???

OpenStudy (anonymous):

\[=-4^2+3^2\]\[=16+9\]\[=?\]

Nnesha (nnesha):

nope you can't solve like that first you have to deal with parenthesis negative sign also have square

Nnesha (nnesha):

@sabby75488 welcome to open study!! please cooperate with us bec u are new here i should tell you that direct answer are not allowed we need you to learn better http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

will it be like this? \[=-4^2-3^2\] @Nnesha

Nnesha (nnesha):

 first take square of -3